OpenStudy (anonymous):

tan theta= -3/4, angle theta in Quadrant IV sin theta= a. 3/5 b. 4/5 c. -4/5 d. -3/5

OpenStudy (anonymous):

For this you are using wolfram alpha ???

OpenStudy (anonymous):

Sorry: \[\large \sec^2(\theta) = 1 + \tan^2 (\theta)\]

OpenStudy (anonymous):

Oh I see. True. theta= 2 pi n + pi

OpenStudy (anonymous):

\[\large \sec^2(\theta) = 1 + \tan^2 (\theta) = 1 + (\frac{-3}{4})^2 \implies \frac{25}{16}\] \[\sec(\theta) = \frac{5}{4}\]

OpenStudy (anonymous):

Cos will be reverse of this: \[\cos(\theta )= \frac{4}{5}\]

OpenStudy (anonymous):

Now use: \[\sin^2 (\theta ) = 1 - \cos^2(\theta) = 1 - \frac{16}{25} = \frac{3}{5}\] But your answer will be negative because \(sin(\theta)\) in 4rt quadrant is negative..

OpenStudy (anonymous):

So: \[\large \color{green}{\sin(\theta) = \frac{-3}{5}}\]

OpenStudy (anonymous):

Oh, O.K. Thank you so much!

OpenStudy (anonymous):

How would I find cot theta and csc theta for this problem?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

\[\cot(\theta)= \frac{1}{\tan(\theta)}\] \[cosec(\theta) = \frac{1}{\sin(\theta)}\]

OpenStudy (anonymous):

-4/3 and -5/3, can you please clarify? @waterineyes I really appreciate your help

OpenStudy (anonymous):

Yes you are absolutely right..